a ball is thrown into the air from a height of 3 feet with an initial velocity of 20 feet per second and is caught at a height of 6 feet. Find the maximum height of the ball and how long it takes to hit the ground.
Answer by stanbon(75887)

a ball is thrown into the air from a height of 3 feet with an initial velocity of 20 feet per second and is caught at a height of 6 feet. Find the maximum height of the ball and how long it takes to hit the ground.
----
h(t) = -16t^2+20t+3
----
Max height occurs when t = -b/(2a) = -20/(2*-16) = 5/8 sec.
----
Height at that time: h(5/8) = 9.25 ft.
----
Ground time:
Solve -16t^2+20t+3 = 0 for "t":
t = [-20 +- sqrt(20^2-4*-16*3)]/-32
---
t = [-20 - sqrt(592)]/-32
t = 1.3853 seconds
============================
Cheers,
Stan H.